# Heath–Jarrow–Morton meet lifted Heston in energy markets for joint historical and implied calibration

### v1

@@ 4.3 Smile shape calibration
We denote the parameters being calibrated by $\mathscr{P}=(\rho,\,c,\,x)$, where the spot-vol correlations $\rho$ are transformed to $\tilde{\rho}$ by (4.22). The optimization problem then reads
@@ 5.1.2 Step 2: Term structure exact calibration correction
In order to avoid jumps of the day-ahead contracts volatility, both functions $g$ and $h$ are smoothed during calibration. At each step of the fixed point algorithm, we update $g$ and $h$ as piece-wise constant functions and then interpolate it using the monotonicity preserving algorithm PCHIP of Fritsch and Butland (1984). Despite the smoother results, the algorithm is more time-consuming due to the numerical optimization involved in the calibration. On the left-hand side of Figure LABEL:fig:calibrated_g below, we plot the results of the term structure calibration of piece-wise constant functions $g$ and $h$, and on the right-hand side, same results with smoothing interpolation. In both cases, the fixed-point algorithm converges in approximately 555510101010 iterations, providing an almost perfect fit for the VS volatilities shown in Figure 14.
@@ 5.2.1 Sample path trajectories errors
As an illustration of the quality of this approximation, we provide in Figure 16 the trajectories of the exact futures contracts $F$ and their approximations $\tilde{F}$, simulated with the same random numbers, for three different futures contracts: Sep 24, Q1 25 and Cal 25. Recall that $F$ is defined in (1.12) with $f(t,T)$ explicitly given by (1.7). The mean $L^{2}$ distance on $[0,\theta]$ is defined as a sample Root Mean Squared Error (RMSE) over $m\in\mathbb{N}^{*}$ sample trajectories by
@@
where $N_{\theta}$ is the number of steps in the daily subdivision of $[0,\theta]$. The estimated distance with $\theta=T_{s}$ between the approximated and exact trajectories estimated over $10^{4}$ simulations equals $0.0602$, $0.0897$ and $0.1898$ for three futures contracts correspondingly. We observed that this error depends both on the interval length $\theta$ as well as on the length of the delivery period. Both dependencies are non-linear, the first one is strictly increasing in $\theta$. As for second one, the error is greater for small (up to one-two months) and for large intervals (two years or more) delivery periods. The worst-case RMSE numerically reaches $0.5\%$ of the initial futures value $F_{0}$.
@@ 5.2.2 Instantaneous correlations
In this section, we assess numerically the impact of the function $g$ on the futures contracts’ instantaneous correlations. Indeed, unlike the function $h$ and the stochastic variance component $V$ which only modify the futures’ volatility term structure but not their correlations, the function ${g}$ also impacts the correlations as it changes the approximated futures contract volatilities $\Sigma_{.}$ defined by (1.17). Indeed, for $i,j\in\{1,\ldots,P\}$, the instantaneous correlation between the approximated futures contracts $\widetilde{F}_{.}^{i}:=\widetilde{F}_{.}^{i}(T_{s}^{i},\,T_{e}^{i})$ and $\widetilde{F}_{.}^{j}:=\widetilde{F}_{.}^{j}(T_{s}^{j},\,T_{e}^{j})$ obtained just after the first calibration step, and implied by the historical futures correlation structure of rolling futures, is defined by
@@
Note that these two values coincide if and only if ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g(T)}$ is constant on $[T_{s}^{i},\,T_{e}^{i}]$ and on $[T_{s}^{j},\,T_{e}^{j}]$, as in this case, $\Sigma_{.}(T_{s}^{i},\,T_{e}^{i})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{g}_{i}}\hat{\Sigma}_{.}(T_{s}^{i},\,T_{e}^{i})$ and $\Sigma_{.}(T_{s}^{j},\,T_{e}^{j})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{g}_{j}}\hat{\Sigma}_{.}(T_{s}^{j},\,T_{e}^{j})$, where $\bar{g}_{i}$ and $\bar{g}_{j}$ denote the values of the function $g$ on the intervals $[T_{s}^{i},\,T_{e}^{i}]$ and $[T_{s}^{j},\,T_{e}^{j}]$ correspondingly.
However, we can show numerically that the difference between $\widetilde{\rho}_{ij}$ and $\widetilde{\rho}_{ij}^{\mathrm{hist}}$ is not significant from a practical point of view even if $g$ is not constant. Namely, we examine this difference for the model calibrated in Section 5.1 with the function $g$ given by Figure LABEL:fig:calibrated_g(b). The figure 17 shows the distances $\|\widetilde{\rho}_{ij}-\widetilde{\rho}_{ij}^{\mathrm{hist}}\|_{\infty}$ for each pair of contracts for the calibration set. In the worst case, the difference is less than 0.1%. Thus, despite the introduction of the function $g$, the correlations in the calibrated model are still very close to the ones given by the historical calibration.
@@ 5.3 Volatility hypercube extrapolation in the calibrated model
To illustrate the volatility term structure, we compute the ATM volatilities of options on daily contracts expiring in the day preceding the day of the delivery, as well as the ATM volatilities of (generated) monthly, quarterly and calendar contracts expiring three days before the delivery period start. The range of considered days cover the days from July 3, 2024 to July 3, 2025. The resulting ATM volatilities of are shown in Figure LABEL:fig:calibrated_vol_ts (a), (b).
We also generate the implied volatility smiles for monthly futures contracts corresponding to the months between August 2024 and December 2026. The options maturity dates are taken equal to the 25-th day of the month preceding the delivery period. The generated smiles are shown in Figure LABEL:fig:calibrated_vol_ts (c).
@@ A.2 Calibration results for TTF market
The monthly contracts surface and the ATM volatilities are shown in Figure LABEL:fig:surf_TTF. Note that the term structure of the ATM volatility is not decreasing for short maturities. This is a consequence of the fact that the implied volatility corresponding to the first contract is $5\%$ lower that the volatilities of the other monthly contracts, possibly due to the seasonality. It is also interesting to notice the the difference between daily and monthly volatility is much smaller than in the power market thanks to the very high correlation between all the futures contracts.
@@ B.2 Linear cone programming
$C_{0}$ :=$:=\left\{u\in\mathbb{R}^{N}\;:\;u_{i}\geq 0,\;i\in\left\{1,\ldots,N\right\}\right\},$ (B.9)
$C_{1}$ :=$:=\left\{(u_{0},u_{1})\in\mathbb{R}\times\mathbb{R}^{\frac{(N+N_{c}+1)(N+N_{c})}{2}}\;:\;u_{0}\geq\|u_{1}\|_{2}\right\},$ (B.10)
$C_{2}$ :=$:=\left\{\mathrm{vec}(u)\;|\;u\in\mathbb{S}_{+}^{N}\right\},$ (B.11)
and where .$\|.\|_{2}$ the standard Euclidean norm, and $\mathrm{vec}(u)$ denotes a symmetric matrix $u$ stored as a vector in column major order.
@@
  • we specify $c:=\left(1,0_{\mathbb{R}^{\frac{(N+N_{c}+1)(N+N_{c})}{2}}}\right)$ so that “$\bar{x}_{0}$ plays the role of $u_{0}$” in the second-order cone $C_{1}$ from (B.10), and we define accordingly the $\tau(a)$-dependent quantity $G^{\lambda}\left(\tau(a)\right)$ using the weights (3.24)–(3.25), and $h$ with the respective target market values from the loss definition (3.22) so as to ensure both the quadratic loss function $J^{\lambda}$ is minimized and $\left(x_{i,j}\right)_{i,j\in I}\in\mathbb{S}_{+}^{N}$, where $I$ denotes the set of indices relative to $\left\{L,S,C_{1}\right\}$ state variables, i.e. $\left(x_{i,j}\right)_{i,j\in I}$ is obtained by withdrawing the columns and rows involving the state variable $C_{2}$ from the matrix $\left(x_{p,k}\right)_{p,k\in\left\{1,\ldots,N+N_{c}\right\}}$,
@@
Without loss of generality, suppose that $[T_{s}^{j},T_{e}^{j}]\subset[T_{s}^{1},T_{e}^{1}]$ for $j\in\{2,\ldots{P_{\mathrm{imp}}}\}$, since the delivery periods non included in $[T_{s}^{1},T_{e}^{1}]$ will not impact the function $h$. Hence, they can be eliminated from the fixed-point iteration algorithm to be calibrated once at the end.
@@
Since we have assumed that the contracts have positive instantaneous correlations, we obtain $(Q_{k})_{ij}>0$ and $(p_{k})_{i}>0$ for all $i,j\in\{2,\ldots,{P_{\mathrm{imp}}}\}$. Moreover, by the linear independence hypothesis on the division, $[T_{s}^{1},T_{e}^{1}]\setminus\cup_{i=2}^{P_{\mathrm{imp}}}[T_{s}^{i},T_{e}^{i}]\not=\varnothing$, so that $r_{k}>0$.
@@ Appendix D Monte Carlo simulation scheme
$U_{t+\Delta t}^{i}=\frac{1}{1+x_{i}\Delta t}\bigl{(}U_{t}^{i}-\lambda V_{t}\Delta t+\nu\sqrt{V_{t}}(B_{t+\Delta t}-B_{t})\bigr{)}\$ (D.2)

### v2

@@ 4.3 Smile shape calibration
We denote the parameters being calibrated by $\mathscr{P}=(\rho,\,c,\,x)$, where the spot-vol correlations $\rho$ are transformed to $\tilde{\rho}$ by (4.22) and (1.6). The optimization problem then reads
@@ 5.1.2 Step 2: Term structure exact calibration correction
In order to avoid jumps of the day-ahead contracts volatility, both functions $g$ and $h$ are smoothed during calibration. At each step of the fixed point algorithm, we update $g$ and $h$ as piece-wise constant functions and then interpolate it using the monotonicity preserving algorithm PCHIP of Fritsch and Butland (1984). Despite the smoother results, the algorithm is more time-consuming due to the numerical optimization involved in the calibration. On the left-hand side of Figure 14(a) below, we plot the results of the term structure calibration of piece-wise constant functions $g$ and $h$, and on the right-hand side, same results with smoothing interpolation. In both cases, the fixed-point algorithm converges in approximately 551010 iterations, providing an almost perfect fit for the VS volatilities shown in Figure 14.
@@ 5.2 A posteriori validation of model approximations
Remark 5.1.
In real energy markets, the smallest possible delivery intervals $\Delta$ are typically 15 minutes, 30 minutes, or one hour. Consequently, the price of any futures contract $F_{t}(T_{s},T_{e})$ can be expressed as the average
$F_{t}(T_{s},T_{e})=\frac{\Delta}{T_{e}-T_{s}}\sum_{i=1}^{(T_{e}-T_{s})/\Delta}f(t,T_{s}+i\Delta),$
of the prices $f(t,T_{s}+i\Delta)$ of all contracts with the smallest granularity $\Delta$ included in $T_{s},T_{e}]$. This implies that the futures curve is inherently discrete, whereas the HJM model provides a continuous approximation. By approximating the contract $F(T_{s},T_{e})$ using formula (LABEL:eq:futures_discrete_approx), we effectively perform an inverse approximation, representing a continuous curve by a discrete sum. For numerical feasibility in the subsequent Monte Carlo simulations, the value $\delta=\frac{1}{365}$, corresponding to one day, was used instead of the finer granularity of one hour or one quarter.
@@ 5.2.1 Sample path trajectories errors
As an illustration of the quality of this approximation, we provide in Figure 16 the trajectories of the exact futures contracts $F$ and their approximations $\tilde{F}$, simulated with the same random numbers, for three different futures contracts: Sep 24, Q1 25 and Cal 25. Recall that $F$ is defined in (1.12) with $f(t,T)$ explicitly given by (1.7). The mean $L^{2}$ distance on $0,\theta]$ is defined as a sample Root Mean Squared Error (RMSE) over $m\in\mathbb{N}^{*}$ sample trajectories by
@@
where $N_{\theta}$ is the number of steps in the daily subdivision of $0,\theta]$. The estimated distance with $\theta=T_{s}$ between the approximated and exact trajectories estimated over $10^{4}$ simulations equals $0.0602$, $0.0897$ and $0.1898$ for three futures contracts correspondingly. We observed that this error depends both on the interval length $\theta$ as well as on the length of the delivery period. Both dependencies are non-linear, the first one is strictly increasing in $\theta$. As for second one, the error is greater for small (up to one-two months) and for large intervals (two years or more) delivery periods. The worst-case RMSE numerically reaches $0.5\%$ of the initial futures value $F_{0}$.
@@ 5.2.2 Instantaneous correlations
In this section, we assess numerically the impact of the function $g$ on the futures contracts’ instantaneous correlations. Indeed, unlike the function $h$ and the stochastic variance component $V$ which only modify the futures’ volatility term structure but not their correlations, the function g{g} also impacts the correlations as it changes the approximated futures contract volatilities $\Sigma_{.}$ defined by (1.19). Indeed, for $i,j\in\{1,\ldots,P\}$, the instantaneous correlation between the approximated futures contracts $\widetilde{F}_{.}^{i}:=\widetilde{F}_{.}^{i}(T_{s}^{i},\,T_{e}^{i})$ and $\widetilde{F}_{.}^{j}:=\widetilde{F}_{.}^{j}(T_{s}^{j},\,T_{e}^{j})$ obtained just after the first calibration step, and implied by the historical futures correlation structure of rolling futures, is defined by
@@
Note that these two values coincide if and only if $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g(T)$} is constant on $T_{s}^{i},\,T_{e}^{i}]$ and on $T_{s}^{j},\,T_{e}^{j}]$, as in this case, $\Sigma_{.}(T_{s}^{i},\,T_{e}^{i})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{g}_{i}}\hat{\Sigma}_{.}(T_{s}^{i},\,T_{e}^{i})$ and $\Sigma_{.}(T_{s}^{j},\,T_{e}^{j})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{g}_{j}}\hat{\Sigma}_{.}(T_{s}^{j},\,T_{e}^{j})$, where $\bar{g}_{i}$ and $\bar{g}_{j}$ denote the values of the function $g$ on the intervals $T_{s}^{i},\,T_{e}^{i}]$ and $T_{s}^{j},\,T_{e}^{j}]$ correspondingly.
However, we can show numerically that the difference between $\widetilde{\rho}_{ij}$ and $\widetilde{\rho}_{ij}^{\mathrm{hist}}$ is not significant from a practical point of view even if $g$ is not constant. Namely, we examine this difference for the model calibrated in Section 5.1 with the function $g$ given by Figure 14(a)(b). The figure 17 shows the distances $\|\widetilde{\rho}_{ij}-\widetilde{\rho}_{ij}^{\mathrm{hist}}\|_{\infty}$ for each pair of contracts for the calibration set. In the worst case, the difference is less than 0.1%. Thus, despite the introduction of the function $g$, the correlations in the calibrated model are still very close to the ones given by the historical calibration.
@@ 5.3 Volatility hypercube extrapolation in the calibrated model
To illustrate the volatility term structure, we compute the ATM volatilities of options on daily contracts expiring in the day preceding the day of the delivery, as well as the ATM volatilities of (generated) monthly, quarterly and calendar contracts expiring three days before the delivery period start. The range of considered days cover the days from July 3, 2024 to July 3, 2025. The resulting ATM volatilities of are shown in Figure 22(a) (a), (b).
We also generate the implied volatility smiles for monthly futures contracts corresponding to the months between August 2024 and December 2026. The options maturity dates are taken equal to the 25-th day of the month preceding the delivery period. The generated smiles are shown in Figure 22(a) (c).
(a) ATM implied volatilities of daily contracts (a) and of monthly, quarterly and calendar contracts (b), and implied volatility smiles for monthly contracts (c) generated by the calibrated model.
@@ A.2 Calibration results for TTF market
The monthly contracts surface and the ATM volatilities are shown in Figure 31(a). Note that the term structure of the ATM volatility is not decreasing for short maturities. This is a consequence of the fact that the implied volatility corresponding to the first contract is $5\%$ lower that the volatilities of the other monthly contracts, possibly due to the seasonality. It is also interesting to notice the the difference between daily and monthly volatility is much smaller than in the power market thanks to the very high correlation between all the futures contracts.
@@
(a) Implied volatility smiles for monthly contracts and implied ATM volatilities of daily and monthly contracts generated by the calibrated model. ((a) and (b) correspondingly).
@@ B.2 Linear cone programming
$C_{0}$ $:=\left\{u\in\mathbb{R}^{N}\;:\;u_{i}\geq 0,\;i\in\left\{1,\ldots,N\right\}\right\},$ (B.9)
$C_{1}$ $:=\left\{(u_{0},u_{1})\in\mathbb{R}\times\mathbb{R}^{\frac{(N+N_{c}+1)(N+N_{c})}{2}}\;:\;u_{0}\geq\|u_{1}\|_{2}\right\},$ (B.10)
$C_{2}$ $:=\left\{\mathrm{vec}(u)\;|\;u\in\mathbb{S}_{+}^{N}\right\},$ (B.11)
and where $\|.\|_{2}$ the standard Euclidean norm, and $\mathrm{vec}(u)$ denotes a symmetric matrix $u$ stored as a vector in column major order.
@@
  • we specify $c:=\left(1,0_{\mathbb{R}^{\frac{(N+N_{c}+1)(N+N_{c})}{2}}}\right)$ so that $\bar{x}_{0}$ plays the role of $u_{0}$” in the second-order cone $C_{1}$ from (B.10), and we define accordingly the $\tau(a)$-dependent quantity $G^{\lambda}\left(\tau(a)\right)$ using the weights (3.24)–(3.25), and $h$ with the respective target market values from the loss definition (3.22) so as to ensure both the quadratic loss function $J^{\lambda}$ is minimized and $\left(x_{i,j}\right)_{i,j\in I}\in\mathbb{S}_{+}^{N}$, where $I$ denotes the set of indices relative to $\left\{L,S,C_{1}\right\}$ state variables, i.e. $\left(x_{i,j}\right)_{i,j\in I}$ is obtained by withdrawing the columns and rows involving the state variable $C_{2}$ from the matrix $\left(x_{p,k}\right)_{p,k\in\left\{1,\ldots,N+N_{c}\right\}}$,
@@
Without loss of generality, suppose that $T_{s}^{j},T_{e}^{j}]\subset[T_{s}^{1},T_{e}^{1}]$ for $j\in\{2,\ldots{P_{\mathrm{imp}}}\}$, since the delivery periods non included in $T_{s}^{1},T_{e}^{1}]$ will not impact the function $h$. Hence, they can be eliminated from the fixed-point iteration algorithm to be calibrated once at the end.
@@
Since we have assumed that the contracts have positive instantaneous correlations, we obtain $(Q_{k})_{ij}>0$ and $(p_{k})_{i}>0$ for all $i,j\in\{2,\ldots,{P_{\mathrm{imp}}}\}$. Moreover, by the linear independence hypothesis on the division, $T_{s}^{1},T_{e}^{1}]\setminus\cup_{i=2}^{P_{\mathrm{imp}}}[T_{s}^{i},T_{e}^{i}]\not$$\varnothing$, so that $r_{k}>0$.
@@ Appendix D Monte Carlo simulation scheme
$U_{t+\Delta t}^{i}=\frac{1}{1+x_{i}\Delta t}\bigl(U_{t}^{i}-\lambda V_{t}\Delta t+\nu\sqrt{V_{t}}(B_{t+\Delta t}-B_{t})\bigr)\\$ (D.2)